the length is 1ft lesss then three times thw width . the perimeter is 22 ft . find the length and width
Answer by bluemockingjay7(41)

perimeter = 22 ft
length = 3x - 1
width = x
3x - 1 + x = 22
2(3x - 1 + x) = 22 (Since 2(L + W)= P)
6x - 2 + 2x = 22
8x - 2 = 22
8x = 22 + 2
8x = 24
8x/8 = 24/8
x = 3
width = 3ft
length = 3(3) - 1 = 9 - 1 = 8ft